# Quantum Molecular Charge-Transfer Model for Multi-step Auger-Meitner Decay Cascade Dynamics

### v1

@@ Abstract
The fragmentation of molecular cations following inner-shell decay processes in molecules containing heavy elements underpins the x-ray damage effects observed in x-ray scattering measurements of biological and chemical materials, as well as in medical applications involving Auger-electron-emitting radionuclides. Traditionally, these processes are modeled using simulations that describe the electronic structure at an atomic level, thereby omitting molecular bonding effects. This work addresses the gap by introducing a novel approach that couples a decay spawning dynamics algorithm with ab initio molecular dynamics simulations to characterize the potential energy surfaces. We apply our method to a model decay cascade following K-shell ionization of IBr and subsequent Kβ fluorescence decay. We examine two competing channels that undergo two decay steps, resulting in ion-pairs with a total +3 charge state. This approach provides a continuous description of the electron transfer dynamics occurring during the multi-step decay cascade and molecular fragmentation, revealing the charge transfer timescale to be approximately 75 fs. Our computed kinetic energies of ion fragments show good agreement with experimental data.
@@ I Introduction
Observing ultrafast molecular dynamics often involves the detection of molecular cation fragments and electrons following inner-shell decay processes, however little is known about the explicit nature of the molecular breakup dynamics[1]. The structural rearrangements following the decay of core-hole states have far reaching implications. They are the fundamental process behind the radiation damage effects caused by structural distortions occurring during x-ray scattering measurements of metallo-protein and clean energy materials[2, 3, 4]. Furthermore, new highly targeted cancer therapies are being developed that exploit Auger-Meitner decay cascades initiated by radionucleotides bonded to small carrier molecules[5, 6]. Therefore, there is demand for the development of simulations that can explicitly track both the ejection of electrons and the associated structural dynamics.
@@
Synchrotron experiments employing x-ray/ion coincidence spectroscopy can discern the charge states and kinetic energies of diverse decay channels subsequent to x-ray-induced molecular breakdown, as noted in [7]. These experiments serve as dependable benchmarks for emerging quantum molecular dynamics methods. We previously employed Monte Carlo/Molecular Dynamics (MC/MD) simulations combined with the classical over-the-barrier (COB) model [8, 9, 7] to study the fragmentation dynamics in IBr molecule induced by x-rays. This approach enables tracking of inner-shell cascades and incorporates electron redistribution and nuclear motion across ground, core-hole, and valence states for all feasible charge configurations. However, it employs the atomic Hartree-Fock-Slater (HFS) method to describe the electronic structure, and the electron transfer COB model promptly fills vacancies in neighboring atoms when the electron binding energy exceeds the Coulomb barrier. Consequently, molecular bonding effects, as well as the quantum effects of electron transfer, are absent from these simulations.
@@
While there has been recent progress in the computation and interpretation of molecular Auger-Meitner decay spectra[10, 11, 12, 13, 14, 15], there have been few developments in quantum molecular dynamics methods for this decay process. Coupling nuclear dynamics to Auger-Meitner decay cascades poses challenges due to the exponential increase in possible final states and potential energy surfaces with system size. Furthermore, the decay of core-hole states in heavy elements, such as bromine and iodine, involves multiple cascade steps, further increasing the complexity of decay dynamics that couples a multitude of potential energy surfaces. Picon introduced a theoretical framework using numerical time-dependant Shrödigner equation simulations[16], applied to in a recent a x-ray pump/probe study[17]. Other studies have focused on vibronically resolving Auger-Meitner decay and coupling it with conformer sampling from dynamics simulation[18, 19, 20, 21]. To our knowledge, no time-resolved studies have yet coupled nuclear dynamics with a multi-step, Auger-Meitner decay dynamics in a heavy-element-containing molecule.
@@
In this work, we provide a quantum molecular description for modelling charge transfer and decay dynamics during the structural damage of molecules following x-ray interactions. To capture the decay dynamics, our approach utilizes a time-dependent ensemble of quantum amplitude trajectories that transfers population by spawning trajectories across potential energy surfaces with increasing charge state. Our approach is inspired by the ab-intio multiple spawning (AIMS) method, which uses a time-dependant set of trajectory basis functions to model the nonadiabatic dynamics associated with the population transfer near conical intersections[22, 23, 24, 25]. To describe the electron reorganization and redistribution in the valence shells and the subsequent charge separation dynamics, we employ ab-inito molecular dynamics simulation. To validate our approach, we use a simplified model to investigate the decay cascade of the IBr molecule, beginning with a K-shell ionization followed by a Kβ fluorescence decay in Br atom. This cascade dynamics has been measured experimentally using an x-ray/ion coincidence. Our calculated results show good agreement with the measured kinetic energy released of the ions and branching ratio of fragmentation channels.
@@
Our paper is organized as followed. In Section II.1, we will describe the decay cascade model for the molecular dynamics of IBr. In Section II.2 and II.3, we present the theoretical treatment and computational details respectively and in Section II.4. Finally, in Section III, we show the calculated results, followed by a summary and outlook.
@@ II Theoretical Method
The experiment was conducted at beamline 7-ID at the Advanced Photon Source, utilizing x-ray/ion coincidence spectroscopy to measure the charge states and energies of Iq+ and Br atomic ions following 1s ionization at the I and Br K-edges of IBr. Detailed experimental procedures are available in our previous work [7], where we presented ion data in coincidence with K emission from both I and Br atoms, along with the mechanisms for ion production derived from the MC/MD method. In the current study, we investigate the decay cascade of the IBr molecule followed by K fluorescence decay in the Br atom. We focus on the dynamics leading to the formation of and . The experimental data were recorded for both 79BrI and 81BrI and similar results were obtained for both isotopes. For our calculations, we uses the 79BrI isotope, for which the the relative yield of to is 1.5 and experimental kinetic energy releases were 12.2 and 11.9 eV, respectively.
@@ II.1 Decay Cascade Model
We use a simplified model to characterize the molecular dynamics. Using MC/MD simulations, we identified two dominant and competing decay channels and molecular states of IBr, starting from a core-excited state with a 3p vacancy in Br atom. Both channels lead to a molecular ion with a total charge state of 3+ via two Auger decays. We index the states based on which step () and channel () they occur in and label the states as . The initial Br 3 hole cation state is considered to be the zeroth step of the zeroth channel and is labelled as the state. In the first decay step, the state decays to the and states in channels 1 and 2 respectively, as shown in Fig. 1. has a vacancy in both the 3d orbital of Br atom and valence orbital, which is delocalized across the molecule. On other hand, has a vacancy in both the 3d and 4s orbital of Br atom. In the second decay step, and decays to and respectively. has a vacancy in 4s orbital of Br atom and two holes in the valence orbital, whereas has two holes in 4s orbital of Br atom and one hole in the valence orbital. The associated Auger-Meitner decay rates () are provided by the MC/MD and HFS simulations respectively and are shown Fig. 1 and discussed in detail elsewhere[8, 7]. We point out the fluorescence decays are not included in the current simplified model
@@ II.2 Coupled AIMD and Decay Spawning Dynamics
This subsection describes the workflow of our coupled decay spawning dynamics and AIMD method and its capability for calculating experimental observables. A single non-interacting potential energy surface is used to approximate all possible electronic configurations for each of the states in the decay channels. In this study, we precalculate all the potential energy surfaces and use them to model the Auger-Meitner decay. Therefore, preliminary ab initio molecular dynamics (AIMD) simulations were performed to produce the non-interacting potential energy surfaces of selected non-ground state cation configurations of the electronic structure. We will discuss the electronic structure calculations in the computational details section. AIMD is a mixed classical-quantum approach which enables the characterization of potential energy surfaces ”on-the-fly”, by classically propagating the nuclei in time with forces determined by an approximate solution to the time-independent electronic Schrödigner equation,
R denotes the nuclear coordinates, and are the electronic Hamiltonian and electronic state wave-functions and , indicates that the electronic structure undergoes self-consistent field optimization at the new nuclear coordinates produced at each time-step. The term represents the nuclear-nuclear repulsion component of the potential energy. Evaluating this equation allows for the propagation of nuclear motion classically by numerically solving the velocity Verlet algorithm [26]. This process yields both the time-dependent nuclear kinetic energies and the atomic charges and predicts the final kinetic and charge states of the ions, which are the relevant experimental observables in x-ray/ion-coincidence experiments.
Characterizing the non-interacting potential energy surfaces using ab initio molecular dynamics (AIMD) facilitates the formation of a total electronic wavefunction for all states involved in the decay cascade model. This wavefunction is parametrically dependent on the nuclear coordinates. Within the Markov approximation, we represent this wavefunction by separating the time-dependent coefficients from their stationary state oscillation in time, as shown below,
@@
To model the Auger-decay induced population transfer between states in the decay cascade, our method utilizes a time-evolving basis of amplitude trajectories that propagate across the predefined AIMD potential energy surfaces. We adopt the spawning nomenclature from AIMS by stating that the state is the parent state with its parent amplitude trajectory . The child amplitude trajectories () for states with will be spawned from , and all states with will have child amplitude trajectories spawned from the states of the corresponding channel. Therefore, there will be only a single amplitude trajectory for , described by . The quantum amplitude trajectories of the states with will be expressed as linear combinations of the child trajectories and are given as follows:
@@
is a user defined parameter that controls the resolution of the wave-packet evolution. If the user was only interested in the final state populations and properties, then would only occur at the final time-step of the simulation. However, in order to capture the charge transfer time of the coupled Auger-Meitner decay and fragmentation dynamics, the selection of should relate to Auger-Meitner decay rates so that the transfer of population between the states is visualized.
In order to relate the results of the decay spawning dynamics to the experimental observables, the properties from the AIMD simulations are coupled to the above EOMs through the decay spawning algorithm described in the next section. This enables population weight averaged properties that describe the overall dynamical behavior of the decay cascade wavepacket by the following equation,
@@ II.3 Computational Details of AIMD Simulations
All AIMD simulations were initiated from the experimental ground state IBr structure with a bond length of 2.469 Å. We excluded the temperature effect and set the initial velocities to zero. A time-step of 0.242 fs was used in the velocity Verlet algorithm, with velocities scaled to maintain constant total energy throughout the AIMD simulations. Eqns 4 and 5 were solved numerically using the fourth-order Runge–Kutta algorithm (RK4). The RK4 method can suffer from numerical population loss for large time-steps. Therefore, to minimize numerical error in the RK4 procedure, the AIMD potential energy curves were fitted with a 1D spline of 100,000 points. The fitted potential energy surfaces were then used in the RK4 simulation with a timestep of 0.00242 fs and the time step and Auger-Meitner decay rates were scaled by 4 orders of magnitude during the RK4 procedure to further reduce the numerical error.
@@ II.4 Decay Spawning Algorithm
Once the potential energy surfaces have been characterized by AIMD, we initiate the spawning dynamics across the surfaces with the initial condition of at , where the spawning decay dynamics begins at of the AIMD simulation. The flow diagram in Fig. 3 and the numbered list below outline how the decay spawning algorithm can be described in three steps. These steps are applicable to any spawning event () for any number of decay cascade steps and channels:
@@
Steps 1-3 can be repeated until the desired property (state population, kinetic energy, atomic charge) has reached a satisfactory convergence. Since our current implementation of this method uses pre-calculated, discretized potential energy surfaces, the initial structures of the individual spawning child trajectories in Step 3 are chosen to have the highest structural similarity to the structure of the parent trajectory at the time of the spawning event. In the case of diatomic IBr, the structural similarity is simply defined as the closest bond length. For polyatomic systems, this metric could be defined by the root mean square deviation (RMSD).
@@ III.1 AIMD
We first discuss the AIMD results associated with the dissociation dynamics of cations involved in the decay cascade. Fig. 4 a) shows the kinetic energy as a function of time. The dication (solid lines) and trication (dashed lines) states converge towards kinetic energies of approximately 3-4 eV and 10-12 eV, respectively, with minor differences between the two channels. Fig. 4 b) and c) show the relative potential energy and bond lengths over time. Since the simulations of these cation states are initiated at the ground state geometry with zero velocities, their initial kinetic energies and bond lengths are the same at . The results indicate that changes in the kinetic energies and bond lengths are more affected by the charge state than by the selected electronic configurations within the charge states. Although there is a strong dependence on the initial potential energy with respect to both the charge state and electronic configuration, the change in potential energy reflects the kinetic energy as the velocity scaling maintains a constant total energy. Fig. 4 d) shows the relative potential energy surfaces of the cation states involved in the decay cascade as a function of bond length. We note that at the final simulation time of 242 fs the potential and kinetic energies have not fully converged as there remains a minor Coulombic interaction between the atomic fragments at the final bond distances.
@@
Using Löwdin electron population analysis, Fig. 5 a) shows the atomic charge populations of the initial cation, di- and tri-cation states. The dashed and dotted lines indicate the atomic populations on Br and I, respectively. In the initial state, the hole is located on the Br 3. However, due to molecular bonding, the atomic charges for Br and I are approximately 0.2 and 0.8, respectively.
@@
For IBr[3d1,V1]+2 dication state, the atomic Br and I atomic charge converge to a 1 and 1 respectively. On the other hand, for IBr[3d1,4s1]+2 dication state, the atomic Br and I charge converge to 0.8 and 1.2 respectively. Similar fractional charges are shown in the dynamics of the trication states. For IBr[4s1,V2]+3, it shows a charge oscillation between the Br and I sites during the first 50 fs before reaching a final atomic Br and I charge state of 2.1 and 0.9. Also, for IBr[4s2,V1]+3, it converges to the atomic Br and I charge of 1.7 and 1.3 respectively.
@@
We interpret the observed non-integer charges on the dissociated atoms as an effect of molecular bonding. Fig. 5c shows the RASSCF molecular hole orbitals in the dication and trication states at both the initial and final structures of the AIMD simulations. It can be seen that the fragmentation of the Br 4s orbital results in hole density on both the Br and I atoms. We attribute this to the mixing of the Br 4s orbital with the I 5s orbital, where the Br 4s contributes to the molecular bonding at the level of theory used in this study. Additionally, the non-integer charges could possibly arise from the finite set of orbitals used in the active spaces and the molecular symmetry restrictions adopted in our simplified model, which neglects other decay processes, including fluorescence. However, it is reasonable to interpret that the two dication states produce ion pairs, whereas IBr[4s2,V1]+3 and IBr[4s,V2]+3 produce and ion pairs, respectively.

### v2

@@ Abstract
The fragmentation of molecular cations following inner-shell decay processes in molecules containing heavy elements underpins the x-ray damage effects observed in x-ray scattering measurements of biological and chemical materials, as well as in medical applications involving Auger-electron-emitting radionuclides. Traditionally, these processes are modeled using simulations that describe the electronic structure at an atomic level, thereby omitting molecular bonding effects. This work addresses the gap by introducing a novel approach that couples a decay spawning dynamics algorithm with ab initio molecular dynamics simulations to characterize ultrafast dynamics on the potential energy surfaces. We apply our method to a model decay cascade following K-shell ionization of IBr and subsequent Kβ fluorescence decay. We examine two competing channels that undergo two decay steps, resulting in ion pairs with a total +3 charge state. This approach provides a continuous description of the electron transfer dynamics occurring during the multi-step decay cascade and molecular fragmentation, revealing the combined inner-shell decay and charge transfer timescale to be approximately 75 fs. Our computed kinetic energies of ion fragments show good agreement with experimental data.
@@ I Introduction
Observing ultrafast molecular dynamics often involves the detection of molecular cation fragments and electrons following inner-shell decay processes, however little is known about the explicit nature of the molecular breakup dynamics[1]. The structural rearrangements following the decay of core-hole states have far-reaching implications. They are the fundamental process behind the radiation damage effects caused by structural distortions occurring during x-ray scattering measurements of metallo-protein and clean energy materials[2, 3, 4]. Also, new highly targeted cancer therapies are being developed that exploit Auger-Meitner decay cascades initiated by radionucleotides bonded to small carrier molecules[5, 6]. Furthermore, Coulomb explosion imaging that uses intense x-ray pulses to create highly charged fragments have been employed to record molecule structures [7, 8, 9]. Therefore, there is demand for the development of simulations that can explicitly track both the ejection of electrons and the associated structural and fragmentation dynamics of x-ray excited molecules.
@@
Synchrotron experiments employing x-ray/ion coincidence spectroscopy can discern the charge states and kinetic energies of diverse decay channels subsequent to x-ray-induced molecular breakdown, as noted in [10]. These experiments serve as reliable benchmarks for emerging quantum molecular dynamics methods. We previously employed Monte Carlo/Molecular Dynamics (MC/MD) simulations combined with the classical over-the-barrier (COB) model [11, 12, 10] to study the fragmentation dynamics in IBr molecule induced by x-rays. This approach enables tracking of inner-shell cascades and incorporates electron redistribution and nuclear motion across the ground, core-hole, and valence states for all feasible charge configurations. However, it employs the atomic Hartree-Fock-Slater (HFS) method to describe the electronic structure, and the electron transfer via the COB model promptly fills vacancies in neighboring atoms when the electron binding energy exceeds the Coulomb barrier. Consequently, molecular bonding effects, as well as the quantum effects of electron transfer, are absent from these simulations.
@@
Although there has been recent progress in the computation and interpretation of molecular Auger-Meitner decay spectra[13, 14, 15, 16, 17, 18], there have been few developments in quantum molecular dynamics methods for this decay process. Coupling nuclear dynamics to Auger-Meitner decay cascades poses challenges because of the exponential increase in possible final states and potential energy surfaces with system size. Furthermore, the decay of core-hole states in heavy elements, such as bromine and iodine, involves multiple cascade steps, further increasing the complexity of decay dynamics that couples a multitude of potential energy surfaces. Picon introduced a theoretical framework using numerical time-dependant Shrödigner equation simulations[19], applied to in a recent x-ray pump/probe study[20]. Other studies have focused on vibronically resolving Auger-Meitner decay and coupling it with conformer sampling from dynamics simulation[21, 22, 23, 24]. To our knowledge, no time-resolved studies have yet coupled nuclear dynamics with a multi-step, Auger-Meitner decay dynamics in a heavy-element-containing molecule.
@@
In this work, we provide a quantum molecular description for modeling charge transfer and decay dynamics during the structural damage of molecules following x-ray interactions. To capture the decay dynamics, our approach utilizes a time-dependent ensemble of quantum amplitude trajectories that transfers population by spawning trajectories across potential energy surfaces with increasing charge state. Our approach is inspired by the ab-intio multiple spawning (AIMS) method, which uses a time-dependant set of trajectory basis functions to model the nonadiabatic dynamics associated with the population transfer near conical intersections[25, 26, 27, 28]. To describe the electron reorganization and redistribution in the valence shells and the subsequent charge separation dynamics, we employ ab-inito molecular dynamics (AIMD)[29] simulation coupled to a novel decay spawning dynamics algorithm. To validate our approach, we use a simplified model to investigate the decay cascade of the IBr molecule, beginning with a K-shell ionization followed by a Kβ fluorescence decay in Br atom. This cascade dynamics has been measured experimentally using an x-ray/ion coincidence. Our calculated results show good agreement with the measured kinetic energy released of the ions and branching ratio of fragmentation channels.
@@
Our paper is organized as follows. In Section II.1, we will describe the decay cascade model for the molecular dynamics of IBr. In Section II.2 and II.3, we present the theoretical treatment and computational details respectively and in Section II.4. Finally, in Section III, we show the calculated results, followed by a summary and outlook.
@@ II Theoretical Method
The experiment was conducted at beamline 7-ID at the Advanced Photon Source, utilizing x-ray/ion coincidence spectroscopy to measure the charge states and energies of Iq+ and Br atomic ions following 1s ionization at the I and Br K-edges of IBr. Detailed experimental procedures are available in our previous work [10], where we presented ion data in coincidence with K emission from both I and Br atoms, along with the mechanisms for ion production derived from the MC/MD method. In the current study, we investigate the decay cascade of the IBr molecule followed by K fluorescence decay in the Br atom. We focus on the dynamics leading to the formation of and . The experimental data were recorded for both 79BrI and 81BrI and similar results were obtained for both isotopes. For our calculations, we uses the 79BrI isotope, for which the the experimental relative yield of to is 1.5 and measured kinetic energy releases were 12.2 and 11.9 eV, respectively.
@@ II.1 Decay Cascade Model
We use a simplified model to characterize the molecular dynamics. Using MC/MD simulations, we identified two dominant and competing decay channels and molecular states of IBr, starting from a core-excited state with a 3p vacancy in Br atom. Both channels lead to a molecular ion with a total charge state of 3+ via two Auger decays. We index the states based on which step () and channel () they occur in and label the states as . The initial Br 3 hole cation state is considered to be the zeroth step of the zeroth channel and is labelled as the state. In the first decay step, the state decays to the and states in channels 1 and 2 respectively, as shown in Fig. 1. has a vacancy in both the 3 orbital of Br atom and valence orbital. On other hand, has a vacancy in both the 3 and orbital. In the second decay step, and decays to and respectively. has a vacancy in orbital and two holes in the valence and orbitals, whereas has two holes in orbital and one hole in the valence orbital. The associated Auger-Meitner decay rates ()shown Fig. 1 are obtained using HFS simulations, which are discussed in detail elsewhere[11, 10]. We point out the possible fluorescence decays, which take at picoseconds or longer, after step 2 are not included in the current simplified model.
@@ II.2 Coupled AIMD and Decay Spawning Dynamics
This subsection describes the workflow of our coupled decay spawning dynamics and AIMD method and its capability for calculating experimental observables. For each of the states in the decay channels, a single non-interacting potential energy surface is used to approximate all possible electronic configurations. In this study, we precalculate all the potential energy surfaces and use them to model the Auger-Meitner decays. Therefore, preliminary AIMD simulations were performed to produce the non-interacting potential energy surfaces of selected non-ground state cation configurations of the electronic structure. We will discuss the electronic structure calculations in the computational details section. AIMD is a mixed classical-quantum approach which enables the characterization of potential energy surfaces “on-the-fly”, by classically propagating the nuclei in time with forces determined by an approximate solution to the time-independent electronic Schrödigner equation,
R denotes the nuclear coordinates, and are the electronic Hamiltonian and electronic state wave-functions, and indicates that the electronic structure undergoes self-consistent field optimization at the new nuclear coordinates produced at each time-step. The term represents the nuclear-nuclear repulsion component of the potential energy. Evaluating this equation allows for the propagation of nuclear motion classically by numerically solving the Velocity Verlet algorithm [30]. This step yields both the time-dependent nuclear kinetic energies and the atomic charges and predicts the final kinetic and charge states of the ions, which are the relevant experimental observables in x-ray/ion-coincidence experiments.
Characterizing the non-interacting potential energy surfaces using AIMD facilitates the formation of a total electronic wavefunction for all states involved in the decay cascade model. This wavefunction is parametrically dependent on the nuclear coordinates. Within the Markov approximation, we represent this wavefunction by separating the time-dependent coefficients from their stationary state oscillation in time, as shown below,
@@
To model the Auger-decay induced population transfer between states, our method utilizes a time-evolving basis of amplitude trajectories that propagate across the predefined AIMD potential energy surfaces. We adopt the spawning nomenclature from AIMS by stating that the state is the parent state with its parent amplitude trajectory . The child amplitude trajectories () for states with will be spawned from , and all states with will have child amplitude trajectories spawned from the states of the corresponding channel. Therefore, there will be only a single amplitude trajectory for , described by . The quantum amplitude trajectories of the states with will be expressed as linear combinations of the child trajectories and are given as follows:
@@
is a user defined parameter that controls the resolution of the wave-packet evolution. If the user was only interested in the final state populations and properties, then would only occur at the final time-step of the simulation. However, in order to capture the charge transfer time of the coupled Auger-Meitner decay and fragmentation dynamics, the selection of should capture Auger-Meitner decay rates so that the transfer of population between the states is visualized.
In order to relate the results of the decay spawning dynamics to the experimental observables, the properties from the AIMD simulations are coupled to the above EOMs through the decay spawning algorithm described in the next section. To compute the population weight averaged properties that describe the overall dynamical behavior of the decay cascade wavepacket, we use the following equation,
@@ II.3 Computational Details of AIMD Simulations
All AIMD simulations were initiated from the experimental ground state IBr structure with a bond length of 2.469 Å. We excluded the temperature effect and set the initial velocities to zero. A time-step of 0.242 fs was used in the velocity Verlet algorithm, with velocities scaled to maintain constant total energy throughout the AIMD simulations. Eqns. 4 and 5 were solved numerically using the fourth-order Runge–Kutta algorithm (RK4). The RK4 method can suffer from numerical population loss for large time-steps. Therefore, to minimize numerical error in the RK4 procedure, the AIMD potential energy curves were fitted with a 1D spline of 100,000 points. The fitted potential energy surfaces were then used in the RK4 simulation with a timestep of 0.00242 fs and the time step and Auger-Meitner decay rates were scaled by 4 orders of magnitude during the RK4 procedure to further reduce the numerical error.
@@ II.4 Decay Spawning Algorithm
Once the potential energy surfaces have been characterized by AIMD, we initiate the spawning dynamics across the surfaces with the initial condition of at , where the spawning decay dynamics begins at of the AIMD simulation. The flow diagram in Fig. 3 shows how the decay spawning algorithm can be executed in three steps. These steps are applicable to any spawning event () for any number of decay cascade steps and channels:
@@
Steps 1-3 can be repeated until the desired properties (state population, kinetic energy, atomic charge) are converged. Since our current implementation of this method uses pre-calculated, discretized potential energy surfaces, the initial structures of the individual spawning child trajectories in Step 3 are chosen to have the highest structural similarity to the structure of the parent trajectory at the time of the spawning event. In the case of diatomic IBr, the structural similarity is simply defined as the closest bond length. For polyatomic systems, this metric could be defined by the root mean square deviation (RMSD).
@@ III.1 AIMD
We first discuss the AIMD results associated with the dissociation dynamics of the cations involved in the decay cascade. Fig. 4 a) shows the kinetic energy as a function of time. The dication (solid lines) and trication (dashed lines) states converge towards kinetic energies of approximately 3-4 eV and 10-12 eV, respectively, with minor differences between the two channels. Fig. 4 b) and c) show the relative potential energy and bond lengths over time. Since the simulations of these cation states are initiated at the ground state geometry with zero velocities, their initial kinetic energies and bond lengths are the same at . The results indicate that changes in the kinetic energies and bond lengths are more affected by the charge state than by the selected electronic configurations within the charge states. Although there is a strong dependence on the initial potential energy with respect to both the charge state and electronic configuration, the change in potential energy reflects the kinetic energy as the velocity scaling maintains a constant total energy. Fig. 4 d) shows the relative potential energy surfaces of the cation states involved in the decay cascade as a function of bond length. We note that at the final simulation time of 242 fs the potential and kinetic energies have not fully converged as there remains a minor Coulombic interaction between the atomic fragments at the final bond distances.
@@
Using Löwdin electron population analysis, Fig. 5 a) shows the atomic charge populations of the initial cation, dication and trication states. The dashed and dotted lines indicate the atomic populations on Br and I, respectively. In the initial state, the hole is located on the Br 3. However, due to molecular bonding, the atomic charges for Br and I are approximately 0.2 and 0.8, respectively.
@@
For IBr[,]+2 dication state, the atomic Br and I atomic charge converge to a 1 and 1 respectively. On the other hand, for IBr[,]+2 dication state, the atomic Br and I charge converge to 0.8 and 1.2 respectively. Similar fractional charges are shown in the dynamics of the trication states. For IBr[,,]+3, it shows a charge oscillation between the Br and I sites during the first 50 fs before reaching a final atomic Br and I charge state of 2.1 and 0.9. Also, for IBr[,]+3, it converges to the atomic Br and I charge of 1.7 and 1.3 respectively.
@@
We interpret the observed non-integer charges on the dissociated atoms as an effect of molecular bonding. Fig. 5c shows the RASSCF molecular hole orbitals in the dication and trication states at both the initial and final structures of the AIMD simulations. It can be seen that the fragmentation of the orbital results in hole density on both the Br and I atoms. This molecular orbital is formed from the Br 4s and I 5s orbitals, which therefore contribute to the molecular bonding at the level of theory used in this study. However, the non-integer charges can possibly be attributed to the finite set of orbitals used in the active spaces, the molecular symmetry restrictions adopted in our simplified model and the neglect of other decay processes, including fluorescence. Nonetheless, it is reasonable to interpret that the two dication states produce ion pairs, whereas IBr[,]+3 and IBr[,,]+3 produce and ion pairs, respectively.